OpenStudy (idku):

A 1250-kg car moves at 22.0 m/s. How much work is required from the engine to increase the car's speed to 36.0 m/s?

rebeccaxhawaii (rebeccaxhawaii):

36/22

OpenStudy (agent0smith):

@rebeccaxhawaii not exactly... Work done = change in kinetic energy \[\large W = KE_f - KE_i\] and KE = 0.5mv^2

OpenStudy (idku):

Yes, so, 0.5(1250)(36)^2-0.5(1250)(22)^2

OpenStudy (idku):

0.5(1250)(14)^2

OpenStudy (idku):

like this?

OpenStudy (agent0smith):

First step yes. Second step no, very much no.

OpenStudy (idku):

Oh, yeah 36^2-22^2\(\ne\)14^2

OpenStudy (agent0smith):

You can not subtract squares like that. 36^2 - 22^2 is absolutely NOT equal to 14^2.

OpenStudy (idku):

I my bad ... that was my worst job factoring:)

rebeccaxhawaii (rebeccaxhawaii):

oops sorry its science related not really my thing but I do remember some

OpenStudy (idku):

tnx

OpenStudy (idku):

That is what I was getting wrong... that math :)

OpenStudy (idku):

Alright, I guess, meeting closed. Thanks for replying Smith!

OpenStudy (agent0smith):

You can factor out all the other stuff though 0.5(1250)(36)^2-0.5(1250)(22)^2 = 0.5(1250)(36^2-22^2)